# Towards Scalable IoT Deployment for Visual Anomaly Detection via Efficient Compression

### v1

@@ Abstract
Visual Anomaly Detection (VAD) is a key task in industrial settings, where minimizing waste and operational costs is essential. Deploying deep learning models within Internet of Things (IoT) environments introduces specific challenges due to the limited computational power and bandwidth of edge devices. This study investigates how to perform VAD effectively under such constraints by leveraging compact and efficient processing strategies. We evaluate several data compression techniques, examining the trade-off between system latency and detection accuracy. Experiments on the MVTec AD benchmark demonstrate that significant compression can be achieved with minimal loss in anomaly detection performance compared to uncompressed data.
@@ 1 Introduction
00footnotetext: This work was partially carried out within the Italian National Center for Sustainable Mobility (MOST) and received funding from NextGenerationEU (Italian NRRP – CN00000023 - D.D. 1033 17/06/2022 - CUP C93C22002750006).
Visual anomaly detection (VAD) is essential in industrial automation for detecting defective parts via image-based inspection. Traditional VAD systems operate in centralized environments with high-performance computing and reliable data acquisition infrastructures [1]. However, with the rise of the Internet of Things (IoT), this paradigm is evolving. IoT enables distributed sensing for real-time asset monitoring across remote environments. Industrial IoT (IIoT) networks connect production sites, supporting flexible monitoring architectures [2, 3]. These changes introduce new constraints for VAD systems, as edge devices have limited resources and bandwidth, requiring the restructuring of VAD pipelines. This has spurred the development of lightweight, communication-efficient frameworks based on unsupervised deep learning and distributed processing [4].
VAD research spans from computer vision to deep learning, as documented in surveys [1]. Studies have explored anomaly detection under resource constraints [5]. Edge-assisted systems reduce latency in industrial monitoring [2], while [3] proposes low-latency VAD for smart cities. Research has also explored edge computing integration [4] and traffic flow analysis [6].
In IIoT, research has focused on communication resource allocation [7] and federated learning for distributed detection [8]. Studies have examined edge-cloud architecture trade-offs [9]. Optimization in VAD includes performance-aware frameworks [10] and adaptive models [11]; whereas, the investigations on effective multi-objective methods aim to reduce latency and memory usage while maintaining accuracy [12].
While a universally accepted taxonomy of VAD methods is lacking, recent studies [13, 14] propose a classification into three main categories: reconstruction-based, synthesizing-based, and embedding-based approaches. Reconstruction-based methods assume anomalies cannot be reconstructed due to their absence in the training data (e.g., visual sensor networks in [15]); they use autoencoders, GANs, diffusion models, and Transformers to reconstruct normal samples, identifying anomalies via pixel-wise or SSIM-based differences. However, overfitting or visually similar anomalies can lead to missed detections. Synthesizing-based methods generate artificial anomalies on normal images, such as CutPaste [16] and DRÆM [17], but struggle to capture real-world defect variability. Finally, feature-based methods use pre-trained models to represent normal data distributions; examples include PaDiM [18], PatchCore [19], and FastFlow [20]. These methods achieve high accuracy but can be sensitive to domain shifts.
@@
  • We investigate two strategies: image compression and lightweight CNN feature extraction with compression on the edge, reducing computational and communication burdens while preserving detection accuracy. Also, we explore compression techniques like WebP for image encoding, and random projection and product quantization for feature encoding, achieving high compression rates with minimal performance loss.
@@ 2 Background and problem formulation
In the sequel, we provide the necessary background on VAD to formalize the optimization problem under analysis.
@@ 2.1 Mathematical preliminaries
We denote the set of CNN feature vectors as =$\mathcal{F}=\{f_{1},\dots,f_{N}\}\subset\mathbb{R}^{d}$, where each $f_{i}$ is obtained by concatenating features from $L$ intermediate layers at the same spatial location. A random subset $\mathcal{F}_{s}\subset\mathcal{F}$ is selected such that $|\mathcal{F}_{s}|=\alpha N$ with $\alpha\in(0,1)$ denoting the sampling ratio.
For an input image $x$, the feature map at layer $\ell$ is denoted by $\phi^{(\ell)}(x)\in\mathbb{R}^{C_{\ell}\times H_{\ell}\times W_{\ell}}$.
The operator
$\text{flatten}:\mathbb{R}^{C\times H\times W}\to\mathbb{R}^{CHW}$
maps a 3D feature tensor to a vector.
The global image embedding is then constructed as
$v=\text{concat}(f^{(1)},f^{(2)},f^{(3)}),$
where concat denotes vector concatenation.
The vector $v\in\mathbb{R}^{d}$ is compressed using product quantization, denoted
@@
yielding the quantized representation $\hat{v}=\text{PQ}(v)$.
2.2 Visual anomaly detection approaches
This study focuses exclusively on feature-based methods, which offer a favorable trade-off between performance and computational cost. Feature-based VAD also enables localized detection and supports compact representations that are amenable to compression. Our previous work [5] demonstrated the viability of feature-based approaches in constrained settings. However, not all tasks can be performed at the edge due to computational limitations. IoT enables the use of more powerful anomaly detection models on the server, where computational and memory constraints are less restrictive. In addition, IoT facilitates the continuous updating of the anomaly detection model, a capability that is challenging to implement in fully edge-based solutions.
@@ 2.3 Problem formulation
We address the task of unsupervised visual anomaly detection (VAD) within an IoT-enabled industrial setting, where resource constraints on edge devices impose limitations on both computational capacity and communication bandwidth. The objective is to design a detection pipeline capable of operating across distributed nodes while minimizing data transmission costs and maintaining detection accuracy.
Let $\mathcal{X}=\{x_{i}\}_{i=1}^{n}$ denote a set of input images acquired by edge devices in an industrial IoT environment. Each image $x_{i}\in\mathbb{R}^{H\times W\times 3}$ must be analyzed for anomalies using an unsupervised visual anomaly detection (VAD) model. Let $\mathcal{E}$ represent the set of edge devices, each subject to constraints on computational capacity and communication bandwidth.
Our goal is to learn a detection function
$f:\mathbb{R}^{H\times W\times 3}\to\mathbb{R},\quad f(x)\in[0,1],$
where $f(x)$ represents an anomaly score, with higher values indicating greater likelihood of abnormality.
The distributed nature of the system imposes two constraints:
  • Computation Constraint: Each edge device $e\in\mathcal{E}$ has limited computational capacity, modeled as an upper bound $C_{e}$ on permissible operations or model size.
@@ 3 Proposed method
To address the problem of unsupervised visual anomaly detection (VAD) in IoT environments with limited computational and communication resources, we propose a distributed architecture consisting of two key components: edge devices and a centralized server. This architecture aims to balance computational load and communication costs while maintaining detection accuracy.
@@ 3.1 Deployment architecture
To address computational heterogeneity and communication constraints in IoT environments, we propose a two-part distributed architecture. The core VAD model is hosted on a centralized server, while edge devices are responsible for either image acquisition and compression or feature extraction. We investigate two separate deployment strategies, each representing an alternative solution tailored to different resources and trade-offs:
  1. 1.
    Image Compression at the Edge: The device acquires an image, compresses it using standard codecs such as WebP, and transmits the compressed image to the server. This approach requires minimal memory overhead and implementation complexity, though encoding may be slower on CPU-bound devices.
  2. 2.
    Feature Extraction at the Edge: The device extracts intermediate features using a lightweight CNN (e.g., MobileNetV2), followed by feature compression. The resulting representation is then transmitted to the server. Although this strategy increases the memory footprint due to the model parameters, it reduces server-side processing and can be advantageous in fog computing scenarios.
@@ 3.2 VAD modeling and training
The VAD model is trained only on normal samples. At inference time, it must produce two outputs: (i) an image-level anomaly score indicating whether the entire image is anomalous or not, and (ii) a pixel-level anomaly map highlighting the spatial regions responsible for the anomaly. The latter enhances interpretability and allows human operators to assess the rationale behind a detection outcome.
@@ 3.3 Feature Compression Techniques
For the scenario involving on-device feature extraction, additional feature compression is applied to reduce communication overhead, considering the following strategies.
Random Sampling (RS): As adopted in several feature-based methods such as PaDiM [18], random sampling of feature maps significantly reduces dimensionality with minimal performance degradation, even at sampling factors of 0.25.
@@
Let $L=\{1,2,3\}$ be the indices of the chosen feature extraction layers of the CNN. For a given input image $x$, let $\phi^{(\ell)}(x)\in\mathbb{R}^{C_{\ell}\times H_{\ell}\times W_{\ell}}$ denote the feature map output at layer $\ell\in L$.
Each feature map is flattened into a vector:
$f^{(\ell)}=\text{flatten}(\phi^{(\ell)}(x))\in\mathbb{R}^{C_{\ell}\cdot H_{\ell}\cdot W_{\ell}}$
@@
$v=\text{concat}(f^{(1)},f^{(2)},f^{(3)})\in\mathbb{R}^{d},\quad\text{with }d=\sum_{\ell\in L}C_{\ell}H_{\ell}W_{\ell}$
This vector $v$ is then compressed via product quantization:
$\hat{v}=\text{PQ}(v),$
and the quantized representation $\hat{v}$ is transmitted to the server for anomaly detection.
@@
Table 1: Anomaly detection performance metrics comparison.
Methods Original Raw Features WebP RS PQ RS+WebP RS+PQ
F1 pxl objects 0.606 0.581 0.568 0.502 0.447 0.386 0.439
@@
overall $\Delta$0.00% -8.20% -2.62% -19.79% -32.76% -40.87% -34.60%
ROC img objects 0.979 0.977 0.976 0.950 0.877 0.877 0.870
@@
overall $\Delta$0.00% -1.43% -0.28% -4.29% -16.94% -14.47% -18.73%
4 Experimental Settings
@@ 4.2 Examined scenarios
In this work, we evaluate various data transmission strategies within an IoT-based visual anomaly detection (VAD) framework. Each scenario reflects a different balance between compression rate, computational cost, and anomaly detection performance. We distinguish between two groups: benchmark scenarios, which are tested end-to-end for detection performance and system resource usage, and scenarios with precomputed metrics, where anomaly detection metrics are taken from previous work [5] and only computational requirements are newly evaluated.
All configurations are based on PatchCore [19] as the anomaly detection method, and all input images from the MVTec AD dataset [23] are uniformly resized to $224\times 224$. This resolution is standard in the VAD literature (e.g., PaDiM [18], CFA [24]) and is compatible with the input sizes of the MobileNetV2 and WideResNet backbones used in our experiments.
4.2.1 Benchmark scenarios.
@@
4.2.2 Scenarios with precomputed metrics.
For the following settings, performance metrics have been sourced from our previous work [5], while timing and memory usage were profiled in this study.
@@ 5.1 Anomaly detection performance
We report the anomaly detection performance at both the image level and pixel level. Image-level performance is measured using the Area Under the ROC Curve (AUC), where the task is to classify whether an image contains any anomaly. Pixel-level performance is evaluated using the F1-score, where the model outputs a heatmap and each predicted pixel is compared against the corresponding ground truth pixel to assess segmentation accuracy.
Table 1 provides a comprehensive summary of the pixel-level (F1) and image-level (ROC AUC) scores, aggregated by objects, textures, and overall performance. In terms of pixel-level F1-score, WebP exhibits only a 2% drop from the baseline, followed by the MobileNetV2 edge feature extractor with an 8% loss. All other methods show significantly larger degradations, ranging from 20% to 40%, indicating poor localization accuracy. On the image-level ROC AUC metric, WebP again performs best, with only a 0.3% reduction from the baseline, while other methods suffer 16% to 20% drops. These results clearly highlight the advantage of using WebP compression for anomaly detection in IoT scenarios.
As shown in Figure 3, in terms of F1-scores, the WebP-based approach is the most consistent compression method in all object and texture categories from the MVTec benchmark. It achieves a performance nearly identical to that of sending the original uncompressed image, demonstrating that it preserves sufficient visual information for effective anomaly localization.
In contrast, the “raw features” method, which uses MobileNetV2 to extract features at the edge and transmits them directly, performs less consistently due to the constrained nature of the edge feature extractor. The WebP method instead transmits the compressed image to the server, where feature extraction is done with the more powerful Wide ResNet-50 backbone, resulting in better segmentation performance.
@@ 5.2 Inference latency
The end-to-end inference latency is evaluated by accounting for edge-side feature extraction and optional compression, transmission over the fixed-bandwidth channel, server-side decoding, optional feature extraction, and anomaly detection using a KNN-based approach as in PatchCore.
@@
Table 2: Measured times for single image inference. All times are in milliseconds.
Method Edge Server
Feat.
extract.
Encode Decode
Feat.
extract.
@@
Original 0.74 109.4
Features 15.0 105.7
RS.25 46.1 102.0
WebP 9.1 0.1 0.6 109.5
PQ 30.0 60.0 0.148 106.8
RS.5+PQ 30.0 17.7 0.102 103.8
RS.5+WebP 30.0 9.3 2.9 103.2
@@
The total times presented in Table 3 assume a fixed channel bandwidth of 100 kB/s, which is a realistic figure for Industrial Internet of Things (IIoT) applications. Although we plan to explore performance under varying bandwidths in future work, the current results already demonstrate substantial efficiency gains, achieving up to 80% reduction in end-to-end inference time, including edge processing, transmission, and server-side computation.
Table 3: Comparison of image processing times across compression strategies. Payload is the size (in kB) sent to the server. Tx Time is the transmission time over a 100 kB/s channel. Total Time includes edge, transmission, and server processing. $\Delta$Time indicates total time change vs. baseline (Original). Lower is better for all metrics.
Method Payload [kB] Tx Time [s] Total Time [s] $\Delta$Time

### v2

@@ Abstract
Visual Anomaly Detection (VAD) is a key task in industrial settings, where minimizing operational costs is essential. Deploying deep learning models within Internet of Things (IoT) environments introduces specific challenges due to limited computational power and bandwidth of edge devices. This study investigates how to perform VAD effectively under such constraints by leveraging compact, efficient processing strategies. We evaluate several data compression techniques, examining the tradeoff between system latency and detection accuracy. Experiments on the MVTec AD benchmark demonstrate that significant compression can be achieved with minimal loss in anomaly detection performance compared to uncompressed data. Current results show up to 80% reduction in end-to-end inference time, including edge processing, transmission, and server computation.
@@ 1 Introduction
Visual anomaly detection (VAD) is essential in industrial automation to detect defective parts by image-based inspection. Traditional VAD systems operate in centralized environments with high performance computing and reliable data acquisition infrastructures [1]. However, with the rise of the Internet of Things (IoT), this paradigm is evolving. IoT enables distributed sensing for real-time asset monitoring in remote environments. Industrial IoT (IIoT) networks connect production sites, supporting flexible monitoring architectures [2, 3]. These changes introduce new constraints for VAD systems, as edge devices have limited resources and bandwidth, requiring the restructuring of VAD pipelines. This has spurred the development of lightweight, communication-efficient frameworks based on unsupervised deep learning and distributed processing [4].
VAD research ranges from computer vision to deep learning, as documented in surveys [1]. Studies have explored anomaly detection under resource constraints [5]. Edge-assisted systems reduce latency in industrial monitoring [2], while [3] proposes low-latency VAD for smart cities. Research has also explored edge computing integration [4], traffic flow analysis [6] and health monitoring systems [7].
In IIoT, research has focused on communication resource allocation [8] and federated learning for distributed detection [9]. Studies have examined edge-cloud architecture trade-offs [10]. Optimization in VAD includes performance-aware frameworks [11] and adaptive models [12]; whereas, investigations on effective multiobjective methods aim to reduce latency and memory usage while maintaining accuracy [13].
While a universally accepted taxonomy of VAD methods is lacking, recent studies [14, 15] propose a classification into three main categories: reconstruction-based, synthesizing-based, and embedding-based approaches. Reconstruction-based methods assume anomalies cannot be reconstructed due to their absence in the training data (e.g., visual sensor networks in [16]; they use autoencoders, GANs, diffusion models, and Transformers to reconstruct normal samples, identifying anomalies via pixel-wise or SSIM-based differences. However, overfitting or visually similar anomalies can lead to missed detections. Synthesizing-based methods generate artificial anomalies on normal images, such as CutPaste [17] and DRÆM [18], but struggle to capture real-world defect variability. Finally, feature-based methods use pre-trained models to represent normal data distributions; examples include PaDiM [19], PatchCore [20], and FastFlow [21]. These methods achieve high accuracy but can be resource-intensive and sensitive to domain shifts.
@@
  • We investigate two strategies: edge image compression and lightweight CNN feature extraction with compression, reducing computational and communication burdens while preserving detection accuracy. Also, we explore compression techniques like WebP for image encoding, and random projection and product quantization for feature encoding, achieving high compression rates with minimal performance loss.
@@ 2 Background and problem formulation
In the sequel we provide the necessary background on VAD to formalize the optimization problem under analysis.
@@ 2.1 Mathematical preliminaries
We denote the set of CNN feature vectors as =$\mathcal{F}=\{f_{1},\dots,f_{N}\}\subset\mathbb{R}^{d}$, where each $f_{i}$ is obtained by concatenating features from $L$ intermediate layers at the same spatial location. A random subset $\mathcal{F}_{s}\subset\mathcal{F}$ is selected such that $|\mathcal{F}_{s}|=\alpha N$ with $\alpha\in(0,1)$ denoting the sampling ratio. For an input image $x$, the feature map at layer $\ell$ is denoted by $\phi^{(\ell)}(x)\in\mathbb{R}^{C_{\ell}\times H_{\ell}\times W_{\ell}}$. The operator $\text{flatten}:\mathbb{R}^{C\times H\times W}\to\mathbb{R}^{CHW}$ maps a 3D feature tensor to a vector. The global image embedding is constructed as
$v=\text{concat}(f^{(1)},\ldots,f^{(m)}),\quad m>1,$
where concat denotes vector concatenation. The vector $v\in\mathbb{R}^{d}$ is compressed using product quantization, denoted
@@
yielding the quantized representation
$\hat{v}=\text{PQ}(v).$ (1)
@@ 2.2 Problem formulation
Let $\mathcal{X}=\{x_{i}\}_{i=1}^{n}$ denote a set of input images acquired by edge devices in an industrial IoT environment. Each image $x_{i}\in\mathbb{R}^{H\times W\times 3}$ must be analyzed for anomalies using an unsupervised VAD model. Also, let $\mathcal{E}$ represent the set of edge devices, each subject to constraints on computational capacity and communication bandwidth. Our goal is to learn a detection function
$p:\mathbb{R}^{H\times W\times 3}\to\mathbb{R},\quad p(x)\in[0,1],$
where $p(x)$ represents an anomaly score, with higher values indicating greater likelihood of abnormality.
The distributed system nature, whose pipeline is composed of two stages $p(x)=f(g(x))$, imposes two constraints:
  • Computation Constraint: Each edge device $e\in\mathcal{E}$ has limited computational capacity, modeled as an upper bound $C_{e}$ on admissible operations/ model size.
@@ 3 Proposed method
In this work we address the task of unsupervised VAD within an IoT-enabled industrial setting, where resource constraints on edge devices impose limitations on both computational capacity and communication bandwidth. In the following, we present effective solutions to the objective defined in Sec. 2.2, aiming to design a detection pipeline that operates across distributed nodes while minimizing data transmission costs and preserving detection accuracy.
@@ 3.1 Deployment architecture
To address computational heterogeneity and communication constraints in IoT environments, we propose a two-part distributed architecture, shown in Fig. 1. The core VAD model is hosted on a centralized server, while edge devices are responsible for either image acquisition and compression, or feature extraction (and optional feature compression). Two deployment strategies are considered:
1) Image Compression at the Edge: The device acquires an image, compresses it using standard codecs such as WebP, and transmits the compressed image to the server. This approach requires minimal memory overhead and implementation complexity, though encoding may be slower on CPU-bound devices.
2) Feature Extraction at the Edge: After image acquisition, the device extracts intermediate features using a lightweight CNN (e.g., MobileNetV2). The resulting representation is then transmitted to the server. Although this strategy increases the memory footprint due to the model parameters, it reduces server-side processing and can be advantageous in fog computing scenarios. In this scenario, the resulting feature vectors are still high-dimensional and would pose a communication bottleneck if transmitted directly. To mitigate this, the pipeline includes a feature compression stage, detailed in Section 3.3, which reduces the transmission size while preserving detection-relevant information.
These strategies are designed as alternatives, enabling flexible deployment depending on the hardware profile of the edge device and the bandwidth constraints of the network. The choice between them directly impacts inference latency, communication load, and overall system scalability.
@@ 3.2 VAD modeling and training
The VAD model is trained solely on normal samples. At inference time, it produces two outputs: (i) an image-level anomaly score and (ii) a pixel-level anomaly map, which enhances interpretability by highlighting anomalous regions. To achieve this, the model adopts feature-based methods for visual anomaly detection, which offer a favorable trade-off between detection performance and computational cost. These methods enable localized detection and compact feature representations suitable for transmission and storage.
Our previous work [5] demonstrated the viability of feature-based approaches in constrained settings. However, not all tasks can be performed at the edge due to computational limitations. IoT enables the deployment of more powerful anomaly detection models on the server, where computational and memory constraints are less restrictive. We thus consider two system variants based on the distribution of computation between edge and server. In the first, feature extraction is performed at the edge using a lightweight CNN, specifically a MobileNetV2 model pretrained on ImageNet. The extracted features are transmitted to the server, where anomaly detection is carried out. In the second, the edge device applies the WebP codec to compress the image, which is then sent to the server. The server performs feature extraction using a WideResNet50 model.
The key distinction in modeling lies in the dimensionality and expressiveness of the features: MobileNetV2 generates lower-dimensional representations than WideResNet50. Since both configurations employ memory bank–based models (e.g., PatchCore or PaDiM) residing on the server, the memory bank derived from MobileNetV2 features is more compact, enabling more efficient inference and greater scalability across many connected devices. Conversely, the use of WideResNet50 combined with WebP compression yields higher anomaly detection performance due to richer feature representations.
The experimental results presented in Sec. 4 will provide a detailed discussion of the architectural trade-offs among bandwidth, scalability, and detection accuracy.
@@ 3.3 Feature Compression Techniques
For the scenario 2) detailed in 3.1, involving on-device feature extraction, additional feature compression is applied to reduce communication overhead. We evaluate several strategies:
Random Sampling. As adopted in several feature-based methods such as PaDiM [19], random sampling of feature maps significantly reduces dimensionality with minimal performance degradation, even at sampling factors of 0.25.
@@
Let $L=\{1,2,3\}$ be the indices of the chosen feature extraction layers of the CNN. For a given input image $x$, let $\phi^{(\ell)}(x)\in\mathbb{R}^{C_{\ell}\times H_{\ell}\times W_{\ell}}$ denote the feature map output at layer $\ell\in L$. Each feature map is flattened into a vector:
$f^{(\ell)}=\text{flatten}(\phi^{(\ell)}(x))\in\mathbb{R}^{C_{\ell}\cdot H_{\ell}\cdot W_{\ell}}.$
@@
$v=\text{concat}(f^{(1)},f^{(2)},f^{(3)})\in\mathbb{R}^{d},\quad\text{with }d=\sum\limits_{\ell\in L}C_{\ell}H_{\ell}W_{\ell}.$
This vector $v$ is then compressed via product quantization in (1) and the quantized representation $\hat{v}$ is transmitted to the server for anomaly detection.
@@
Table 1: Comparison of anomaly detection performance metrics.
Metric Class Original Raw Features WebP RS PQ RS+WebP RS+PQ
F1 (pxl) Objects 0.606 0.581 0.568 0.502 0.447 0.386 0.439
@@
$\Delta$ Overall (%) 0.00% $-8.20$% $-2.62$% $-19.79$% $-32.76$% $-40.87$% $-34.60$%
ROC (img) Objects 0.979 0.977 0.976 0.950 0.877 0.877 0.870
@@
$\Delta$ Overall (%) 0.00% $-1.43$% $-0.28$% $-4.29$% $-16.94$% $-14.47$% $-18.73$%
4 Experimental results
@@ 4.2 Examined scenarios
In this work, we evaluate various data transmission strategies within an IoT-based visual anomaly detection (VAD) framework. Each scenario reflects a different balance between compression rate, computational cost, and anomaly detection performance.
All configurations are based on PatchCore [20] as the anomaly detection method, and all input images from the MVTec AD dataset [24] are uniformly resized to $224\times 224$. This resolution is standard in the VAD literature e.g., PaDiM [19], CFA [25] and is compatible with the input sizes of the MobileNetV2 and WideResNet backbones used in our experiments.
@@ 4.3 Anomaly detection performance
We evaluate anomaly detection performance at both the image and pixel levels. Image-level accuracy is assessed using the Area Under the ROC Curve (ROC AUC), measuring the ability to classify whether an image contains any anomaly. Pixel-level performance is measured using the F1-score, which evaluates the accuracy of the predicted anomaly heatmap against ground-truth segmentation masks.
Table 1 summarizes the image-level and pixel-level results, aggregated by object and texture categories, as well as overall performance. On the pixel-level metric, the WebP-based approach shows only a 2% drop in F1-score compared to the uncompressed baseline, while the MobileNetV2 edge feature extractor results in an 8% drop. All remaining methods degrade significantly, with losses between 20% and 40%, indicating a substantial reduction in localization accuracy.
For image-level anomaly detection, WebP again performs best among the lightweight methods, with only a 0.3% reduction in ROC AUC relative to the baseline. Other methods exhibit larger losses ranging from 16% to 20%, highlighting their reduced classification reliability.
Overall, the WebP compression strategy preserves sufficient visual information to maintain high segmentation and classification performance across all MVTec categories, demonstrating its effectiveness and robustness. In contrast, the edge-feature approach using MobileNetV2 yields less consistent results, primarily due to the limited representational capacity of the lightweight feature extractor. These findings, summarized by Fig. 3, confirm that deferring feature extraction to the server —after high-quality compression— offers a better trade-off between communication cost and detection accuracy in resource-constrained IoT deployments. However, edge-side feature extraction remains advantageous in scenarios where server-side scalability and load balancing are critical, as it eliminates the need for server-side feature extraction and reduces the size of the memory bank, as discussed in Sec. 3.2 and previous work [5].
@@ 4.4 Inference latency
The end-to-end inference latency is evaluated by accounting for edge-side feature extraction and optional compression, transmission over the fixed-bandwidth channel, server-side decoding, optional feature extraction, and anomaly detection using a KNN-based approach as in PatchCore. A detailed breakdown of total times for each scenario is reported in Table 3.
@@
Table 2: Inference times per image (ms).
Method Edge Server Tot.
Feature
extracttion
Encoding Decoding
Feature
extraction
@@
Original 0.74 109.4 110.1
Features 15.0 105.7 120.7
RS.25 46.1 102.0 148.1
WebP 9.1 0.1 0.6 109.5 119.3
PQ 30.0 60.0 0.15 106.8 197.0
RS.5+PQ 30.0 17.7 0.10 103.8 151.6
RS.5+WebP 30.0 9.3 2.9 103.2 145.4
@@
The results assume a fixed transmission bandwidth of 100 kB/s, which is a realistic figure for Industrial IoT (IIoT) applications. Total time accounts for edge processing, transmission delay (Tx Time), and server-side computation. Payload refers to the data volume sent to the server. Although we plan to explore varying bandwidths in future work, the current results already demonstrate substantial efficiency gains, achieving up to 80% reduction in end-to-end inference time.
Table 3: Image processing times. Payload is the size of the data sent to the server. Tx Time is the transmission time over a 100 kB/s channel. Total Time includes edge, transmission, and server processing. $\Delta$Time indicates total time change vs. baseline (Original). Lower is better for all metrics.
Method Payload Tx Time Total Time $\Delta$Time
[kB] [s] [s] [vs. Orig]